# Isometric immersions into three-dimensional unimodular metric Lie groups

### v1

@@ 2.2. A unified model in coordinates
If $c_{3}=0$, then $D\times\mathbb{R}=\mathbb{R}^{2}\times\mathbb{R}$ is globally isometric to a semidirect product with a left-invariant metric, and hence we are done (see Example 2.8). Next, by changing all the signs of the $c_{i}$, we can assume that $c_{3}<0$ and $c_{1},c_{2}\geq 0$. On the one hand, if $c_{1}>0$ and $c_{2}>0$, then we have seen in Example 2.6 that $D\times\mathbb{R}$ is globally isometric to $\widetilde{\mathrm{SL}}_{2}(\mathbb{R})$, so we are done too. On the other hand, if $c_{1}=c_{2}=0$, then $D\times\mathbb{R}$ is $\mathbb{R}^{3}$ with the global orthonormal frame $E_{1}=\partial_{x}-\frac{c_{3}}{2}y\partial_{z}$, $E_{2}=\partial_{y}+\frac{c_{3}}{2}x\partial_{z}$ and $E_{3}=\partial_{z}$. This is the standard model for $\mathrm{Nil}_{3}(-\frac{c_{3}}{2})=\mathbb{E}(0,-\frac{c_{3}}{2})$, which is also global. We are left with the case $c_{1}=0$ and $c_{2}>0$ ($c_{1}>0$ and $c_{2}=0$ can be reasoned similarly). Note that this means $D\times\mathbb{R}$ is the slab of $\mathbb{R}^{3}$ defined by the inequality $01+\frac{c_{2}c_{3}}{4}x^{2}>$ 0. The map $F:D\times\mathbb{R}\to\mathbb{R}^{3}$ given by
@@ 3.1. The fundamental data
We will revisit briefly some basic facts about semi-Riemannian immersions, all of which are discussed by O’Neill [25, Ch. 3 and 4], towards Gauss and Codazzi equations. The second fundamental form $\sigma$ of the immersion is the normal component of the connection, that is, we have
¯$\overline{\nabla}_{X}Y=\nabla_{X}Y+\sigma(X,Y),\quad\text{for all }X,Y\in\mathfrak{X}(\Sigma),$
@@
[…]
@@
Items ((((i))))((((ii)))) and ((((iii)))) have been already discussed, so we will focus on ((((iv)))) and ((((v)))). To this end, we will write ¯$\overline{\nabla}_{X}E_{i}$ in two different ways. On the one hand,
¯$\overline{\nabla}_{X}E_{i}=\overline{\nabla}_{X}(T_{i}+\hat{\epsilon}_{3}\nu_{i}N)$ $=\nabla_{X}T_{i}+\sigma(X,T_{i})+\hat{\epsilon}_{3}\langle\nabla\nu_{i},X\rangle N+\hat{\epsilon}_{3}\nu_{i}\overline{\nabla}_{X}N$ (3.9)
@@
On the other hand, ¯$\overline{\nabla}_{X}E_{i}=\sum_{j=1}^{3}\epsilon_{j}\langle X,E_{j}\rangle\overline{\nabla}_{E_{j}}E_{i}$. Expanding this last identity by means of (2.4) and comparing with (3.9), the tangent components give the equations for $\nabla_{X}T_{i}$ whilst the normal components give the equations for $\nabla\nu_{i}$. ∎
These elements $S,J,T_{1},T_{2},T_{3},\nu_{1},\nu_{2},\nu_{3}$ will be called the fundamental data of the immersion. The rotation $J$ has been included here to highlight the role of the orientation and how it must be compatible with the definition of the normal; equivalently, one can assume that $\Sigma$ is oriented a priori. All the conditions in Proposition 3.1 on the fundamental data will be called the fundamental equations. We must say they have a high degree of redundancy, e.g., one can differentiate ((((iii)))) in the direction of a vector field $X\in\mathfrak{X}(\Sigma)$ to obtain some relations between between ((((iv)))) and ((((v)))). This is similar to the case of $\mathbb{E}(\kappa,\tau)$-spaces as pointed out by Daniel [6, Rmk. 4.2]. However, our main goal is to prove that either $T_{1},T_{2},T_{3}$ or $\nu_{1},\nu_{2},\nu_{3}$ are enough to decide if the immersion exists and to recover it just in terms of some subset of the above compatibility equations.
@@
Given an isometric immersion $\phi:\Sigma\to G$ and $\Psi\in\operatorname{Iso}(G)$, we get another isometric immersion $\widetilde{\phi}=\Psi\circ\phi:\Sigma\to G$. If $\Psi$ is a left-translation, then $\phi$ and $\widetilde{\phi}$ clearly have the same fundamental data. If $\Psi\in\operatorname{Stab}_{p}$ preserves the orientation and is not the identity, then $\Psi$ is an axial symmetry about one of the directions $\{E_{1},E_{2},E_{3}\}$ (see [23, Prop. 2.21]). In our Cartan model, $\Psi(x,y,z)$ is equal to $(x,-y,-z)$, $(-x,y,-z)$ or $(-x,-y,z)$, so $\Psi$ preserves the orientation and the change of fundamental data from $\phi$ to $\widetilde{\phi}$ is straightforward. For instance, the symmetry about $E_{1}$ preserves $S$ and $J$ while it changes $(T_{1},T_{2},T_{3})\mapsto(T_{1},-T_{2},-T_{3})$ and $(\nu_{1},\nu_{2},\nu_{3})\mapsto(\nu_{1},-\nu_{2},-\nu_{3})$. The other rotations about $E_{2}$ and $E_{3}$ are similar. This observation will simplify some forthcoming discussions.
@@ 3.2. The change-of-frame equation
The exterior derivatives of all these $1$-forms are well known (e.g., the proof in [7, Prop. 2.4] can be easily adapted to the semi-Riemannian case):
@@
¯$\overline{\nabla}_{e_{k}}e_{\beta}=\sum_{\delta=1}^{3}\left(\mathrm{d}M^{\delta}_{\beta}(e_{k})E_{\delta}+M^{\delta}_{\beta}\overline{\nabla}_{e_{k}}E_{\delta}\right)=\sum_{\zeta=1}^{3}\mathrm{d}M^{\zeta}_{\beta}(e_{k})E_{\zeta}+\!\!\!\!\sum_{\zeta,\gamma,\delta=1}^{3}\!\!M^{\delta}_{\beta}M^{\gamma}_{k}\overline{\Gamma}{}^{\zeta}_{\gamma\delta}E_{\zeta}.$
We can also express ¯$\overline{\nabla}_{e_{k}}e_{\beta}=\sum_{\delta=1}^{3}\omega^{\delta}_{\beta}(e_{k})e_{\delta}$ by means of the curvature forms (3.10), whence a change of basis yields ¯$\overline{\nabla}_{e_{k}}e_{\beta}=\sum_{\zeta,\delta=1}^{3}\omega^{\delta}_{\beta}(e_{k})M^{\zeta}_{\delta}E_{\zeta}$. By matching coefficients in the above two expressions, we reach
@@
We will proceed as in the proof of [6, Prop. 4.11] by showing the integrability of the kernel of $\Lambda=B^{-1}\mathrm{d}\chi-M\omega$, where $\chi=(x,y,z)^{t}$ is the position vector. Note that $\Lambda$ is an array containing three $1$-forms whose kernel is a $3$-dimensional distribution in the $5$-dimensional manifold $\Sigma\times D\times\mathbb{R}$. However, our argument to prove integrability will be different from [6].
@@
The existence of a local immersion around $p_{0}\in\Sigma$ follows as in the last paragraph of [6, Prop. 4.11]. The global existence of the immersion follows from the simple connectedness of $\Sigma$ by a standard continuation argument (see [7, Thm. 3.3]). ∎
@@ 3.3. Recovering the immersion from the tangent projections
Next result shows that the shape operator is also uniquely determined and the compatibility equations ((((iv)))) are enough to recover the immersion.
@@
Let $G$ be the semi-Riemannian unimodular metric Lie group of diagonalizable type with structure constants $c_{1},c_{2},c_{3}$ and signs $\epsilon_{1},\epsilon_{2},\epsilon_{3}$. Let $\Sigma$ be a simply connected semi-Riemannian surface with a positive orthonormal frame $\{e_{1},e_{2}\}$ with signature $\epsilon_{1},\epsilon_{2}$, and assume that $\Sigma$ has three tangent vectors fields $T_{1},T_{2},T_{3}$ such that $\sum_{\alpha=1}^{3}\epsilon_{\alpha}\langle T_{\alpha},e_{i}\rangle\langle T_{\alpha},e_{j}\rangle=\epsilon_{i}\delta^{j}_{i}$ for all $i,j\in\{1,2\}$. Consider the smooth functions $\nu_{1},\nu_{2},\nu_{3}$ given by (3.29). If the compatibility equations (italic-(((iv)))) hold true for some field of self-adjoint operators $S$, then there is a unique isometric immersion $\phi:\Sigma\to G$ for which $T_{1},T_{2},T_{3}$ are the tangent components of the left-invariant frame in $G$.
@@
Define $M^{\alpha}_{j}=\epsilon_{\alpha}\langle T_{\alpha},e_{\beta}\rangle$ and $M^{\alpha}_{3}=\epsilon_{\alpha}\nu_{\alpha}$ for $\alpha\in\{1,2,3\}$ and $k\in\{1,2\}$. The condition $\sum_{\alpha=1}^{3}\epsilon_{\alpha}\langle T_{\alpha},e_{i}\rangle\langle T_{\alpha},e_{j}\rangle=\epsilon_{i}\delta^{j}_{i}$ means that the first two columns of $M$ are orthogonal with norms $\epsilon_{1}$ and $\epsilon_{2}$, respectively, which guarantees that the third column has norm $\epsilon_{3}$ and $M\in\operatorname{SO}^{\epsilon}_{3}(\mathbb{R})$. This is equivalent to satisfying the algebraic conditions ((((iii)))). We also define $\omega^{3}_{j}(e_{k})=\epsilon_{3}\langle Se_{k},e_{j}\rangle$ so that $\omega^{3}_{1}\wedge\omega^{1}+\omega^{3}_{2}\wedge\omega^{2}=0$ is clearly equivalent to the fact that $S$ is self-adjoint.
In view of Remark 3.5, we just need to check that the first two columns of $\mathrm{d}M$ and $M\Omega+ML(M)$ agree, which is equivalent to check that the six identities in (3.20) are satisfied. We will show how to deal with the first one since the others are very similar. Since we are assuming ((((iv)))), we can expand the differential of $M^{1}_{1}$ as
@@
Finally, notice that $S$ must coincide with the shape operator of the immersion (provided that it exists). Moreover, we can employ ((((iv)))) and (3.3) to get
@@ 3.4. Recovering the immersion from the angle functions
We shall assume for a moment that $\Sigma$ is actually immersed in $G$ with angles $\nu_{1},\nu_{2},\nu_{3}$ to gain some intuition. From the compatibility equations ((((iii)))) and ((((v)))), along with the fact that $JS+SJ=2H\epsilon_{3}J$, it is easy to check that
@@
It might seem that (3.38) uniquely determines $T_{\alpha}$ only in terms of $\nu_{1},\nu_{2},\nu_{3}$ and $J$, because $H^{2}$ can be retrieved from (3.37). Unfortunately, the sign of $H$ is undefined and we will see that different signs potentially lead to non-congruent immersions. However, the existence of such immersions will depend on whether or not (3.38) (for $H$ or $-H$) satisfies the hypotheses of Theorem 3.6 when the shape operator $S$ is necessarily given by (3.30). In the next result, we show that the algebraic relations ((((iii)))) and the symmetry of $S$ turn out to be automatically satisfied so the only nontrivial restrictions are the compatibility equations ((((iv)))).
@@
    For each such choice of $H\in\mathcal{C}^{\infty}(\Sigma)$, the existence of the immersion only depends on whether or not $T_{1},T_{2},T_{3}$ and $S$ defined by (3.38) and (3.30), respectively, satisfy the compatibility equations (italic-(((iv)))).
@@
If the immersion exists, then $H$ must be a smooth function satisfying (3.37), and conversely, for each such $H$, the immersion exists if and only if $T_{1},T_{2},T_{3}$ and $S$ satisfy Theorem 3.6. On the one hand, from (3.38) and item (b) of Lemma 3.7, we get the algebraic relations $\langle T_{\alpha},T_{\beta}\rangle=\psi^{-1}\langle X_{\alpha},X_{\beta}\rangle=\epsilon_{\alpha}\delta^{\alpha}_{\beta}-\epsilon_{3}\nu_{\alpha}\nu_{\beta}$. Therefore, the three rows of the matrix $M$ in (3.28) form an orthonormal frame, and hence the columns also form an orthonormal frame, that is, $\sum_{\alpha=1}^{3}\epsilon_{\alpha}\langle T_{\alpha},e_{i}\rangle\langle T_{\alpha},e_{j}\rangle=\epsilon_{i}\delta^{j}_{i}$ for all $i,j\in\{1,2\}$ on a positive orthonormal frame $\{e_{1},e_{2}\}$ with signature $\epsilon_{1},\epsilon_{2}$. On the other hand, the condition for the symmetry of $S$ in Theorem 3.6 can be written as $\sum_{\alpha=1}^{3}\epsilon_{\alpha}\langle\nabla\nu_{\alpha},JT_{\alpha}\rangle=-\epsilon_{1}\epsilon_{2}\epsilon_{3}\zeta$, which by (3.38) can be in turn transformed into $\frac{\zeta}{\psi}\sum_{\alpha=1}^{3}\epsilon_{\alpha}\langle J\nabla\nu_{\alpha},X_{\alpha}\rangle+\frac{2H}{\psi}\sum_{\alpha=1}^{3}\epsilon_{\alpha}\langle\nabla\nu_{\alpha},X_{\alpha}\rangle=\zeta$, and this last identity follows readily from item (c) of Lemma 3.7. All in all, Theorem 3.6 says that the immersion exists if and only if ((((iv)))) are fulfilled, and uniqueness is guaranteed up to left-translations, that is, items (a) and (b) in the statement are proved, except for the last assertion in item (a).
We will see next that the compatibility equations ((((v)))) are also satisfied automatically (this simplifies the forthcoming arguments). Observe that ((((v)))) holds true if and only if it holds true when multiplied by any $T_{\alpha}$. As for the first equation (the other two can be proved similarly), this means we have to show that
@@
for all $\alpha\in\{1,2,3\}$. As in the proof of Lemma 3.7, this computation is burdensome but again it follows from the algebraic relations for $T_{\alpha}$ we have already proved (it is very useful to check first that (3.3) holds just by these algebraic relations) plus the spherical constraints given by (3.34) and (3.35). Knowing that ((((v)))) is true, we can go back to the last sentence in item (a), and compute
$\operatorname{tr}(S)$ =$\stackrel{{\scriptstyle(\star)}}{{=}}\sum_{\alpha=1}^{3}\epsilon_{\alpha}\langle ST_{\alpha},T_{\alpha}\rangle=-\sum_{\alpha=1}^{3}\epsilon_{\alpha}\langle\nabla\nu_{\alpha},T_{\alpha}\rangle$
@@
where we have used (3.38) and item (c) of Lemma 3.7. We would like to remark that the equality ()($\star$)( ⋆ ) is true for any operator and follows from elementary linear algebra along with the algebraic relations.
@@
Finally, we will assume that $H$ is positive in order to prove item (d). Let $T_{1},T_{2},T_{3},S$ be the data of the associated immersion. The first observation is that the equation for $\nabla_{X}T_{\alpha}$ in ((((iv)))), when multiplied by $T_{\alpha}$ is automatically fulfilled. For instance, when $\alpha=1$, we get from ((((v)))) and the symmetry of $S$ that
@@
which agrees with ((((iv)))) innerly multiplied by $T_{1}$. Therefore, the most interesting information in $\nabla_{X}T_{\alpha}$ is its component in the direction $JT_{\alpha}$. Furthermore, the other immersion corresponding to the opposite sign $\widetilde{H}=-H$ will exist if and only if after expanding both sides of ((((iv)))) multiplied by $JT_{\alpha}$, we end up with an even function of $H$. For simplicity, write $T_{\alpha}=aX_{\alpha}+bJX_{\alpha}$ with $a=\frac{\epsilon_{1}\epsilon_{2}\epsilon_{3}\zeta}{\psi}$ and $b=\frac{2H\epsilon_{3}}{\psi}$ as in (3.38), so being even in $H$ amounts to being even in $b$. On the one hand, the left-hand side of ((((iv)))) gives
@@
where $E(b)$ collects all terms which are even in $b$ (notice that the gradient of the logarithm ignores the sign of $H$). On the other hand, the right-hand side of ((((iv)))) multiplied by $JT_{\alpha}$ can be expanded by means of ((((iii)))) and ((((v)))) plus the symmetry of $S$, and then simplified by means of the spherical constraints (3.34) and (3.35) as in the proof of Lemma 3.7. After onerous but simple computations, we reach

### v2

@@ 2.2. A unified model in coordinates
If $c_{3}=0$, then $D\times\mathbb{R}=\mathbb{R}^{2}\times\mathbb{R}$ is globally isometric to a semidirect product with a left-invariant metric, and hence we are done (see Example 2.8). Next, by changing all the signs of the $c_{i}$, we can assume that $c_{3}<0$ and $c_{1},c_{2}\geq 0$. On the one hand, if $c_{1}>0$ and $c_{2}>0$, then we have seen in Example 2.6 that $D\times\mathbb{R}$ is globally isometric to $\widetilde{\mathrm{SL}}_{2}(\mathbb{R})$, so we are done too. On the other hand, if $c_{1}=c_{2}=0$, then $D\times\mathbb{R}$ is $\mathbb{R}^{3}$ with the global orthonormal frame $E_{1}=\partial_{x}-\frac{c_{3}}{2}y\partial_{z}$, $E_{2}=\partial_{y}+\frac{c_{3}}{2}x\partial_{z}$ and $E_{3}=\partial_{z}$. This is the standard model for $\mathrm{Nil}_{3}(-\frac{c_{3}}{2})=\mathbb{E}(0,-\frac{c_{3}}{2})$, which is also global. We are left with the case $c_{1}=0$ and $c_{2}>0$ ($c_{1}>0$ and $c_{2}=0$ can be reasoned similarly). Note that this means $D\times\mathbb{R}$ is the slab of $\mathbb{R}^{3}$ defined by the inequality $1+\frac{c_{2}c_{3}}{4}x^{2}>0$. The map $F:D\times\mathbb{R}\to\mathbb{R}^{3}$ given by
@@ 3.1. The fundamental data
We will revisit briefly some basic facts about semi-Riemannian immersions, all of which are discussed by O’Neill [26, Ch. 3 and 4], towards Gauss and Codazzi equations. The second fundamental form $\sigma$ of the immersion is the normal component of the connection, that is, we have
$\overline{\nabla}_{X}Y=\nabla_{X}Y+\sigma(X,Y),\quad\text{for all }X,Y\in\mathfrak{X}(\Sigma),$
@@
This linear map is self-adjoint with respect to $\langle\cdot,\cdot\rangle$, but not necessarily diagonalizable if $\Sigma$ is Lorentzian. However, we can define the extrinsic and the mean curvature of the immersion, as the determinant and a half of the trace of $S$, respectively:
@@
Items ((i)), ((ii)) and ((iii)) have been already discussed, so we will focus on ((iv)) and ((v)). To this end, we will write $\overline{\nabla}_{X}E_{i}$ in two different ways. On the one hand,
$\overline{\nabla}_{X}E_{i}=\overline{\nabla}_{X}(T_{i}+\hat{\epsilon}_{3}\nu_{i}N)$ $=\nabla_{X}T_{i}+\sigma(X,T_{i})+\hat{\epsilon}_{3}\langle\nabla\nu_{i},X\rangle N+\hat{\epsilon}_{3}\nu_{i}\overline{\nabla}_{X}N$ (3.9)
@@
On the other hand, $\overline{\nabla}_{X}E_{i}=\sum_{j=1}^{3}\epsilon_{j}\langle X,E_{j}\rangle\overline{\nabla}_{E_{j}}E_{i}$. Expanding this last identity by means of (2.4) and comparing with (3.9), the tangent components give the equations for $\nabla_{X}T_{i}$ whilst the normal components give the equations for $\nabla\nu_{i}$. ∎
These elements $S,J,T_{1},T_{2},T_{3},\nu_{1},\nu_{2},\nu_{3}$ will be called the fundamental data of the immersion. The rotation $J$ has been included here to highlight the role of the orientation and how it must be compatible with the definition of the normal; equivalently, one can assume that $\Sigma$ is oriented a priori. All the conditions in Proposition 3.1 on the fundamental data will be called the fundamental equations. We must say they have a high degree of redundancy, e.g., one can differentiate ((iii)) in the direction of a vector field $X\in\mathfrak{X}(\Sigma)$ to obtain some relations between between ((iv)) and ((v)). This is similar to the case of $\mathbb{E}(\kappa,\tau)$-spaces as pointed out by Daniel [7, Rmk. 4.2]. However, our main goal is to prove that either $T_{1},T_{2},T_{3}$ or $\nu_{1},\nu_{2},\nu_{3}$ are enough to decide if the immersion exists and to recover it just in terms of some subset of the above compatibility equations.
@@
Given an isometric immersion $\phi:\Sigma\to G$ and $\Psi\in\operatorname{Iso}(G)$, we get another isometric immersion $\widetilde{\phi}=\Psi\circ\phi:\Sigma\to G$. If $\Psi$ is a left-translation, then $\phi$ and $\widetilde{\phi}$ clearly have the same fundamental data. If $\Psi\in\operatorname{Stab}_{p}$ preserves the orientation and is not the identity, then $\Psi$ is an axial symmetry about one of the directions $\{E_{1},E_{2},E_{3}\}$ (see [24, Prop. 2.21]). In our Cartan model, $\Psi(x,y,z)$ is equal to (x,y,z)(x,-y,-z), (x,y,z)(-x,y,-z) or (x,y,z)(-x,-y,z), so $\Psi$ preserves the orientation and the change of fundamental data from $\phi$ to $\widetilde{\phi}$ is straightforward. For instance, the symmetry about $E_{1}$ preserves $S$ and $J$ while it changes $(T_{1},T_{2},T_{3})\mapsto(T_{1},-T_{2},-T_{3})$ and $(\nu_{1},\nu_{2},\nu_{3})\mapsto(\nu_{1},-\nu_{2},-\nu_{3})$. The other rotations about $E_{2}$ and $E_{3}$ are similar. This observation will simplify some forthcoming discussions.
@@ 3.2. The change-of-frame equation
The exterior derivatives of all these $1$-forms are well known (e.g., the proof in [8, Prop. 2.4] can be easily adapted to the semi-Riemannian case):
@@
$\overline{\nabla}_{e_{k}}e_{\beta}=\sum_{\delta=1}^{3}\left(\mathrm{d}M^{\delta}_{\beta}(e_{k})E_{\delta}+M^{\delta}_{\beta}\overline{\nabla}_{e_{k}}E_{\delta}\right)=\sum_{\zeta=1}^{3}\mathrm{d}M^{\zeta}_{\beta}(e_{k})E_{\zeta}+\!\!\!\!\sum_{\zeta,\gamma,\delta=1}^{3}\!\!M^{\delta}_{\beta}M^{\gamma}_{k}\overline{\Gamma}{}^{\zeta}_{\gamma\delta}E_{\zeta}.$
We can also express $\overline{\nabla}_{e_{k}}e_{\beta}=\sum_{\delta=1}^{3}\omega^{\delta}_{\beta}(e_{k})e_{\delta}$ by means of the curvature forms (3.10), whence a change of basis yields $\overline{\nabla}_{e_{k}}e_{\beta}=\sum_{\zeta,\delta=1}^{3}\omega^{\delta}_{\beta}(e_{k})M^{\zeta}_{\delta}E_{\zeta}$. By matching coefficients in the above two expressions, we reach
@@
We will proceed as in the proof of [7, Prop. 4.11] by showing the integrability of the kernel of $\Lambda=B^{-1}\mathrm{d}\chi-M\omega$, where $\chi=(x,y,z)^{t}$ is the position vector. Note that $\Lambda$ is an array containing three $1$-forms whose kernel is a $3$-dimensional distribution in the $5$-dimensional manifold $\Sigma\times D\times\mathbb{R}$. However, our argument to prove integrability will be different from [7].
@@
The existence of a local immersion around $p_{0}\in\Sigma$ follows as in the last paragraph of [7, Prop. 4.11]. The global existence of the immersion follows from the simple connectedness of $\Sigma$ by a standard continuation argument (see [8, Thm. 3.3]). ∎
@@ 3.3. Recovering the immersion from the tangent projections
Next result shows that the shape operator is also uniquely determined and the compatibility equations ((iv)) are enough to recover the immersion.
@@
Let $G$ be the semi-Riemannian unimodular metric Lie group of diagonalizable type with structure constants $c_{1},c_{2},c_{3}$ and signs $\epsilon_{1},\epsilon_{2},\epsilon_{3}$. Let $\Sigma$ be a simply connected semi-Riemannian surface with a positive orthonormal frame $\{e_{1},e_{2}\}$ with signature $\epsilon_{1},\epsilon_{2}$, and assume that $\Sigma$ has three tangent vectors fields $T_{1},T_{2},T_{3}$ such that $\sum_{\alpha=1}^{3}\epsilon_{\alpha}\langle T_{\alpha},e_{i}\rangle\langle T_{\alpha},e_{j}\rangle=\epsilon_{i}\delta^{j}_{i}$ for all $i,j\in\{1,2\}$. Consider the smooth functions $\nu_{1},\nu_{2},\nu_{3}$ given by (3.29). If the compatibility equations ((iv)) hold true for some field of self-adjoint operators $S$, then there is a unique isometric immersion $\phi:\Sigma\to G$ for which $T_{1},T_{2},T_{3}$ are the tangent components of the left-invariant frame in $G$.
@@
Define $M^{\alpha}_{j}=\epsilon_{\alpha}\langle T_{\alpha},e_{\beta}\rangle$ and $M^{\alpha}_{3}=\epsilon_{\alpha}\nu_{\alpha}$ for $\alpha\in\{1,2,3\}$ and $k\in\{1,2\}$. The condition $\sum_{\alpha=1}^{3}\epsilon_{\alpha}\langle T_{\alpha},e_{i}\rangle\langle T_{\alpha},e_{j}\rangle=\epsilon_{i}\delta^{j}_{i}$ means that the first two columns of $M$ are orthogonal with norms $\epsilon_{1}$ and $\epsilon_{2}$, respectively, which guarantees that the third column has norm $\epsilon_{3}$ and $M\in\operatorname{SO}^{\epsilon}_{3}(\mathbb{R})$. This is equivalent to satisfying the algebraic conditions ((iii)). We also define $\omega^{3}_{j}(e_{k})=\epsilon_{3}\langle Se_{k},e_{j}\rangle$ so that $\omega^{3}_{1}\wedge\omega^{1}+\omega^{3}_{2}\wedge\omega^{2}=0$ is clearly equivalent to the fact that $S$ is self-adjoint.
In view of Remark 3.5, we just need to check that the first two columns of $\mathrm{d}M$ and $M\Omega+ML(M)$ agree, which is equivalent to check that the six identities in (3.20) are satisfied. We will show how to deal with the first one since the others are very similar. Since we are assuming ((iv)), we can expand the differential of $M^{1}_{1}$ as
@@
Finally, notice that $S$ must coincide with the shape operator of the immersion (provided that it exists). Moreover, we can employ ((iv)) and (3.3) to get
@@ 3.4. Recovering the immersion from the angle functions
We shall assume for a moment that $\Sigma$ is actually immersed in $G$ with angles $\nu_{1},\nu_{2},\nu_{3}$ to gain some intuition. From the compatibility equations ((iii)) and ((v)), along with the fact that $JS+SJ=2H\epsilon_{3}J$, it is easy to check that
@@
It might seem that (3.38) uniquely determines $T_{\alpha}$ only in terms of $\nu_{1},\nu_{2},\nu_{3}$ and $J$, because $H^{2}$ can be retrieved from (3.37). Unfortunately, the sign of $H$ is undefined and we will see that different signs potentially lead to non-congruent immersions. However, the existence of such immersions will depend on whether or not (3.38) (for $H$ or H-H) satisfies the hypotheses of Theorem 3.6 when the shape operator $S$ is necessarily given by (3.30). In the next result, we show that the algebraic relations ((iii)) and the symmetry of $S$ turn out to be automatically satisfied so the only nontrivial restrictions are the compatibility equations ((iv)).
@@
    For each such choice of $H\in\mathcal{C}^{\infty}(\Sigma)$, the existence of the immersion only depends on whether or not $T_{1},T_{2},T_{3}$ and $S$ defined by (3.38) and (3.30), respectively, satisfy the compatibility equations ((iv)).
@@
If the immersion exists, then $H$ must be a smooth function satisfying (3.37), and conversely, for each such $H$, the immersion exists if and only if $T_{1},T_{2},T_{3}$ and $S$ satisfy Theorem 3.6. On the one hand, from (3.38) and item (b) of Lemma 3.7, we get the algebraic relations $\langle T_{\alpha},T_{\beta}\rangle=\psi^{-1}\langle X_{\alpha},X_{\beta}\rangle=\epsilon_{\alpha}\delta^{\alpha}_{\beta}-\epsilon_{3}\nu_{\alpha}\nu_{\beta}$. Therefore, the three rows of the matrix $M$ in (3.28) form an orthonormal frame, and hence the columns also form an orthonormal frame, that is, $\sum_{\alpha=1}^{3}\epsilon_{\alpha}\langle T_{\alpha},e_{i}\rangle\langle T_{\alpha},e_{j}\rangle=\epsilon_{i}\delta^{j}_{i}$ for all $i,j\in\{1,2\}$ on a positive orthonormal frame $\{e_{1},e_{2}\}$ with signature $\epsilon_{1},\epsilon_{2}$. On the other hand, the condition for the symmetry of $S$ in Theorem 3.6 can be written as $\sum_{\alpha=1}^{3}\epsilon_{\alpha}\langle\nabla\nu_{\alpha},JT_{\alpha}\rangle=-\epsilon_{1}\epsilon_{2}\epsilon_{3}\zeta$, which by (3.38) can be in turn transformed into $\frac{\zeta}{\psi}\sum_{\alpha=1}^{3}\epsilon_{\alpha}\langle J\nabla\nu_{\alpha},X_{\alpha}\rangle+\frac{2H}{\psi}\sum_{\alpha=1}^{3}\epsilon_{\alpha}\langle\nabla\nu_{\alpha},X_{\alpha}\rangle=\zeta$, and this last identity follows readily from item (c) of Lemma 3.7. All in all, Theorem 3.6 says that the immersion exists if and only if ((iv)) are fulfilled, and uniqueness is guaranteed up to left-translations, that is, items (a) and (b) in the statement are proved, except for the last assertion in item (a).
We will see next that the compatibility equations ((v)) are also satisfied automatically (this simplifies the forthcoming arguments). Observe that ((v)) holds true if and only if it holds true when multiplied by any $T_{\alpha}$. As for the first equation (the other two can be proved similarly), this means we have to show that
@@
for all $\alpha\in\{1,2,3\}$. As in the proof of Lemma 3.7, this computation is burdensome but again it follows from the algebraic relations for $T_{\alpha}$ we have already proved (it is very useful to check first that (3.3) holds just by these algebraic relations) plus the spherical constraints given by (3.34) and (3.35). Knowing that ((v)) is true, we can go back to the last sentence in item (a), and compute
$\operatorname{tr}(S)$ $\stackrel{{\scriptstyle(\star)}}{{=}}\sum_{\alpha=1}^{3}\epsilon_{\alpha}\langle ST_{\alpha},T_{\alpha}\rangle=-\sum_{\alpha=1}^{3}\epsilon_{\alpha}\langle\nabla\nu_{\alpha},T_{\alpha}\rangle$
@@
where we have used (3.38) and item (c) of Lemma 3.7. We would like to remark that the equality $(\star)$ is true for any operator and follows from elementary linear algebra along with the algebraic relations.
@@
Finally, we will assume that $H$ is positive in order to prove item (d). Let $T_{1},T_{2},T_{3},S$ be the data of the associated immersion. The first observation is that the equation for $\nabla_{X}T_{\alpha}$ in ((iv)), when multiplied by $T_{\alpha}$ is automatically fulfilled. For instance, when $\alpha=1$, we get from ((v)) and the symmetry of $S$ that
@@
which agrees with ((iv)) innerly multiplied by $T_{1}$. Therefore, the most interesting information in $\nabla_{X}T_{\alpha}$ is its component in the direction $JT_{\alpha}$. Furthermore, the other immersion corresponding to the opposite sign $\widetilde{H}=-H$ will exist if and only if after expanding both sides of ((iv)) multiplied by $JT_{\alpha}$, we end up with an even function of $H$. For simplicity, write $T_{\alpha}=aX_{\alpha}+bJX_{\alpha}$ with $a=\frac{\epsilon_{1}\epsilon_{2}\epsilon_{3}\zeta}{\psi}$ and $b=\frac{2H\epsilon_{3}}{\psi}$ as in (3.38), so being even in $H$ amounts to being even in $b$. On the one hand, the left-hand side of ((iv)) gives
@@
where $E(b)$ collects all terms which are even in $b$ (notice that the gradient of the logarithm ignores the sign of $H$). On the other hand, the right-hand side of ((iv)) multiplied by $JT_{\alpha}$ can be expanded by means of ((iii)) and ((v)) plus the symmetry of $S$, and then simplified by means of the spherical constraints (3.34) and (3.35) as in the proof of Lemma 3.7. After onerous but simple computations, we reach